OpenStudy (anonymous):

Sarah wants to deposit $2,125 into savings accounts at three different banks: Bank of the US, 7/4 Bank, and Catch Bank. She will deposit four times as much into Catch Bank as 7/4 Bank because they offer a higher interest rate. She also expects the Bank of the US deposit to be 25% of the total of the other two deposits. How much will Sheila deposit into the Bank of the US savings account?

OpenStudy (konradzuse):

I'm not sure 100% how to do this, but we need to look at it in this way. X is the amount we want to find for 7/4 bank. 7/4 bank = x Catch = 4x US Bank is 0.25(25%) of the TOTAL of the others. So I would say 0.25(5x)

OpenStudy (konradzuse):

So it seems like we are looking at x + 4x + (0.25(5x)) = 2125?

OpenStudy (konradzuse):

can you try and solve that for me please?

OpenStudy (anonymous):

ok wait...

OpenStudy (anonymous):

340?

OpenStudy (konradzuse):

you are correct that x = 340, now what are the amounts per each bank?

OpenStudy (anonymous):

i dont know...

OpenStudy (konradzuse):

As I said above. X is the amount we want to find for 7/4 bank. 7/4 bank = x Catch = 4x US Bank is 0.25(25%) of the TOTAL of the others. So I would say 0.25(5x)

OpenStudy (anonymous):

i honestly i don't know what you want me to do? what im so post to plug in?

OpenStudy (konradzuse):

yessir.

OpenStudy (anonymous):

425?

OpenStudy (konradzuse):

There are 3 banks... We found x to be 340 so that's 7/4 bank, now what is 4x, and 0.25(5x) = 425 yes.

OpenStudy (anonymous):

4X=1360

OpenStudy (anonymous):

4(340)=1360

OpenStudy (konradzuse):

Correct, and so what is 340 + 1360 + 425?

OpenStudy (anonymous):

2125=D

OpenStudy (konradzuse):

There we go! understand it better now? Math is a mind game, conquer it and you can take on the world!

OpenStudy (anonymous):

so wait what do i do now? whats the answer?

OpenStudy (konradzuse):

"How much will Sheila deposit into the Bank of the US savings account?"

OpenStudy (anonymous):

425=D?

OpenStudy (konradzuse):

yessir I believe that is the correct choice.

OpenStudy (anonymous):

thanks so much your the best

OpenStudy (konradzuse):

Np, tag me @KonradZuse if you have anymore questions.

OpenStudy (anonymous):

Jeffrey earned a gross income of $78,350 last year. He made $2,489.78 in student loan interest deductions, donated $4,294 to his favorite charities and paid $3,571.32 in home mortgage interest. Jeffrey claims a standard deduction of $11,400 for himself and his non-working spouse. If their exemption is $7,300, what is their taxable income?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

that's right

OpenStudy (konradzuse):

Thanks sir :).